# Fast Generation of Feasible Trajectories in Direct Optimal Control

### v1

@@ V-B Fixed-Time P2P Motion of Unstable System
In this test problem, FP-DDP is compared against a Direct Single Shooting (DSS) implementation using the globalization of FP-DDP and a Direct Multiple Shooting (DMS) implementation using the steering penalty method in [4]. The example that is discussed originates from [5] and is adopted from [2]. The dynamics of the considered system are given by
$\dot{x}_{1}=x_{2}+u\left(\zeta+(1-\zeta)x_{2}\right),$
$\dot{x}_{2}=x_{1}+u\left(\zeta-4(1-\zeta)x_{2}\right),$
@@
Fig. 1 shows the convergence of the different methods. Looking at the iteration numbers reveals that FP-DDP takes 5 iterations to converge to a feasible solution. Interestingly, always full steps are taken. DSS takes 18 iterations until the optimal solution is found. Since the underlying system is highly unstable the open-loop forward simulation in DSS yields trajectories that are far from the solution. This results in smaller steps for DSS and therefore slower convergence. Looking at DMS reveals that the choice of the penalty parameter highly influences the convergence speed. Setting the initial penalty parameter $\sigma=1$ as in [4] achieves convergence in 94 iterations. The Maratos effect seems to be taking place whereas taking an initial penalty parameter of $0.01$ yields convergence within 5 iterations. This demonstrates the dependence of merit-function-based algorithms on their penalty parameter whereas FP-DDP does not require this parameter choice.
V-C Free-Time P2P Motion of Cart Pendulum with Obstacle
@@
In this section, a free-time cart pendulum problem with obstacle avoidance is discussed. The system dynamics and parameters are given in [19]. The states $x=(p,v,\theta,\omega)^{\top}\in\mathbb{R}^{4}$ consist of position, velocity, pole angle, and pole angular velocity. The control (force) inputs are given by $u=F\in\mathbb{R}$. The free-time $T$ is modeled as an additional state using time scaling. The task is to move from an initial upward pole position corresponding to $\bar{x}_{1}=(0,0,0,0)^{\top}$ to the final upward position $\bar{x}_{N+1}=(5.0,0,0,0)^{\top}$. In addition, some bound constraints on the states, i.e., $\underline{x}=(-1.0,-2.0,-\frac{\pi}{4},-0.5)$, $\bar{x}=(6.0,2.0,\frac{\pi}{4},0.5)$, $\underline{x}\leq x_{i}\leq\bar{x}$ with $i=2,\ldots,N$, as well as $|u_{i}|\leq 5.0$ for $i=1,\ldots,N$ are imposed. During the motion, the endpoint of the pole has to avoid a circular obstacle with its center at $(\tilde{p},0.9)^{\top}$, $\tilde{p}\in\mathbb{R}$ and having a radius of $0.3$. The horizontal position of the center of the obstacle is chosen from $100$ linearly spaced values between $0.7$ and $4.3$, i.e., in total, $100$ problems are solved.
@@
FP-DDP is compared against the state-of-the-art solver IPOPT [21] which solves the same problem formulation as FP-DPP and against scipy.optimize.minimize using a Truncated-Newton method [17], [20] and solving the unconstrained NLP of SEQUOIA which is given in Table I. IPOPT is used through its CasADi interface which provides the derivative information. The derivatives for the unconstrained NLP were manually built in CasADi and then passed to scipy.optimize.minimize. As Hessian, the GGN matrix was used. IPOPT used exact Hessian information. The solvers were initialized with a constant trajectory at the initial position. The time is initialized with the constant value $T=5.0$.
@@
To compare the different algorithms, the Dolan and Moré performance profiles [8] were used. As the performance metric the number of Hessian evaluations was chosen. Fig. 2 shows the results of the different solvers. The vertical axis $\rho_{s}(\tau)$ indicates what fraction of all test problems were solved. The horizontal axis $\tau$ indicates that the solver was not more than $\tau$ times worse than the best solver. If the performance of an algorithm is better relative to another algorithm, this is indicated by higher values of $\rho_{s}(\tau)$.
Figure 2: Performance plot of Hessian evaluations for FP-DDP, IPOPT, and truncated Newton Scipy.
@@
It can be seen that FP-DDP outperforms the other algorithms and is in almost 55% of the cases the algorithm with the least amount of Hessian evaluations until convergence is achieved. IPOPT fails once in the restoration phase. Apart from that all solvers are robust and can solve the test problems, but overall FP-DDP converges the fastest. This indicates the strong performance of the proposed algorithm. An additional advantage of FP-DDP is that it can be efficiently implemented using high-performance linear algebra packages such as BLASFEO [11] yielding fast execution times.
VI Conclusion
@@
This paper introduced an efficient algorithm for finding feasible trajectories in direct optimal control based on DDP and an Armijo-type line search. Theoretical convergence guarantees were derived and the efficiency of the algorithm was demonstrated against state-of-the-art methods. Possible future research areas include an efficient C implementation of the algorithm and its embedding into an OCP NLP solver.

### v2

@@ 5.2 Fixed-Time P2P Motion of Unstable System
In this test problem, FP-DDP is compared against a Direct Single Shooting (DSS) implementation using the globalization of FP-DDP and a Direct Multiple Shooting (DMS) implementation using the steering penalty method described in [4]. The example originates from [5] and is adopted from [2]. The dynamics are given by
$\dot{x}_{1}\!=\!x_{2}\!+\!u\left(\zeta+(1\shortminus\zeta)x_{2}\right),\>\dot{x}_{2}\!=\!x_{1}\!+\!u\left(\zeta\shortminus 4(1\shortminus\zeta)x_{2}\right),$
@@
Fig. 1 shows the convergence of the different methods. FP-DDP takes 5 iterations to converge to a feasible solution. Interestingly, always full steps are taken. DSS takes 18 iterations until the optimal solution is found. Since the underlying system is highly unstable, the open-loop forward simulation in DSS yields trajectories that are far from the solution. This results in smaller steps for DSS and therefore slower convergence. Looking at DMS reveals that the choice of the penalty parameter highly influences the convergence speed. Setting the initial penalty parameter $\sigma=1$ as in [4] results in convergence in 94 iterations. The Maratos effect seems to occur. For $\sigma=0.01$, the algorithm convergences within 5 iterations. This demonstrates the dependence of merit-function-based algorithms on their penalty parameter whereas FP-DDP does not require this parameter choice.
5.3 Free-Time P2P Motion of Cart Pendulum
@@
This section discusses a free-time cart pendulum problem with obstacle avoidance. The system dynamics and parameters are given in [20]. The states $x=(p,v,\theta,\omega)^{\top}\in\mathbb{R}^{4}$ consist of position, velocity, pole angle, and pole angular velocity. The control (force) input is given by $u=F\in\mathbb{R}$. The free-time $T$ is modeled as an additional state using time scaling. The task is to move from an initial upward pole position corresponding to $\bar{x}_{1}=(0,0,0,0)^{\top}$ to the final upward position $\bar{x}_{N+1}=(5.0,0,0,0)^{\top}$. In addition, some bound constraints $\underline{x}\leq x_{i}\leq\bar{x}$ with $i=2,\ldots,N$ and $\underline{x}=(-1.0,-2.0,-\frac{\pi}{4},-0.5)$, $\bar{x}=(6.0,2.0,\frac{\pi}{4},0.5)$, as well as $|u_{i}|\leq 5.0$ for $i=1,\ldots,N$ are imposed. During the motion, the endpoint of the pole has to avoid a circular obstacle with its center at $(\tilde{p},0.9)^{\top}$, $\tilde{p}\in\mathbb{R}$ and with a radius of $0.3$. The horizontal position of the center of the obstacle is chosen from $100$ linearly spaced values between $0.7$ and $4.3$, i.e., in total, $100$ problems are solved.
@@
FP-DDP is compared against IPOPT [22], which solves the same problem formulation as FP-DPP, and against scipy.optimize.minimize using a Truncated-Newton method [17], [21] and solving the unconstrained NLP of SEQUOIA which is given in Table 1. IPOPT is called through its CasADi interface which provides the derivative information and enables code generation. Note that code generation increases execution speed, but does not alter the algorithm. The derivatives for the unconstrained NLP as well as the GGN Hessian are manually built in CasADi and passed to scipy. IPOPT uses exact Hessian information. The solvers were initialized with a constant trajectory at the initial position and constant value $T=5.0$.
@@
To compare the different algorithms, the Dolan and Moré performance profiles [7] are used. A problem is solved correctly if the final iterate is dynamically feasible and has an objective function value smaller than $10^{-8}$. Otherwise, it is a failure. As performance metrics, we regard the wall time and the number of Hessian evaluations. Fig. 2 shows the results of the different solvers. The vertical axis indicates the number of test problems that were solved. The horizontal axis on the left shows the absolute wall time and on the right side the ratio of Hessian evaluations needed to converge relative to the best solver, i.e., the solver did not need more than its Hessian evaluation ratio times the number of Hessian evaluations of the best solver until convergence. Higher values on the vertical axis in plot (b) indicate better performance of an algorithm relative to another algorithm.
Figure 2: Performance plot of wall time (left) and Hessian evaluations ratio with respect to the best solver (right).
@@
It can be seen that FP-DDP outperforms the other algorithms and is in almost 55% of the cases the algorithm with the least amount of Hessian evaluations until convergence is achieved. IPOPT fails once in the restoration phase. Apart from that all solvers are robust and converge towards feasible trajectories. Overall, the wall time of FP-DDP implemented within acados is an order of magnitude lower than the code generated IPOPT. Due to the slow run time, scipy was excluded from the wall time plot. This demonstrates the strong performance of FP-DDP.
6 Conclusion
@@
This paper introduced an efficient algorithm for finding feasible trajectories in direct optimal control based on DDP and an Armijo-type line search. Theoretical convergence guarantees were derived and the efficiency of the algorithm was demonstrated against state-of-the-art methods. Possible future research areas include the embedding of FP-DDP into an OCP NLP solver.